Serial numbers for a product are to be made using 3 letters followed by 2numbers. If the letters are to be taken from the first 5 letters of the alphabet with repeats possible and the numbers are taken from the digits 0 through 9 with no repeats, how many serial numbers can be generated?
Answer by stanbon(75887)

Serial numbers for a product are to be made using 3 letters followed by 2numbers. If the letters are to be taken from the first 5 letters of the alphabet with repeats possible and the numbers are taken from the digits 0 through 9 with no repeats, how many serial numbers can be generated?
---------
Pattern:: LLLNN
Ans: 5*5*5*10*10 = 125*100 = 12,500
-------------
Cheers,
Stan H.
-----------